OpenStudy (anonymous):

What is the value of the expression when a = 4? 2a^3

OpenStudy (anonymous):

hey its jango

OpenStudy (jango_in_dtown):

2a^3 put 4 in place of a and you will get 2*4^3=2*64=??

OpenStudy (anonymous):

what does that equal

OpenStudy (anonymous):

@shadybubba

OpenStudy (anonymous):

2(4)^3

OpenStudy (anonymous):

I noticed @jango_IN_DTOWN

OpenStudy (anonymous):

Nvm

OpenStudy (jango_in_dtown):

lol

OpenStudy (anonymous):

512

OpenStudy (anonymous):

Hope it helped

OpenStudy (anonymous):

2*4 = 8 8*8*8 = 512

OpenStudy (jango_in_dtown):

its 128

OpenStudy (anonymous):

XD how on earth did you get 128?

OpenStudy (anonymous):

Oh ... There aren't parentheses there XD Nvm, he is correct

OpenStudy (anonymous):

It is indeed 128